Lifespan: Assume the average life-span of those born in the U.S. is 78.2 years with a standard deviation of 16 years. The distribution is not normal (it is skewed left). The good people at Live-Longer-USA (fictitious) claim that their regiment of acorns and exercise results in longer life. So far, 40 people on this program have died and the mean age-of-death was 84.7 years.

(a) Calculate the probability that a random sample of 40 people from the general population would have a mean age-of-death greater than 84.7 years. Round your answer to 4 decimal places.

(b) Which statement best describes the situation for those in the Live Longer program? Since the probability of getting a sample of 40 people with a mean age-of-death greater than those in the Live Longer program is so small, this suggests that people enrolled in the program do actually live longer on average. This provides solid evidence that acorns and exercise have nothing to do with age-of-death. This provides solid evidence that acorns and exercise cause people to live longer.

(c) Why could we use the central limit theorem here despite the parent population being skewed? Because the sample size is greater than 20. Because skewed-left is almost normal. Because the sample size is greater than 30. Because the sample size is less than 100.
Answer #1

a)

= P(Z > 2.57)

= 1 - P(Z < 2.57)

= 1 - 0.9949

= 0.0051

B) This provides solid evidence that acorns and exercise cause people to live longer.

C) Because the sample size is greater than 30.